# Optimized HDBSCAN clustering for reconstructing the merger history of the Milky Way: applications and limitations

### v1

@@ Abstract
Clustering algorithms can help reconstruct the assembly history of the Milky Way by identifying groups of stars sharing similar properties in a kinematical or chemical abundance space. However, although being promising tools, their efficiency has not yet been fully tested in a realistic cosmological framework.
We investigate the effectiveness of the HDBSCAN clustering algorithm in the recovery of the progenitors of Milky Way-type galaxies, using several systems from the Auriga suite of simulations. We develop a methodology aimed at improving the efficiency of the algorithm and avoiding fragmentation: First, we use a 12-dimensional feature space including a range of chemodynamical properties and stellar ages; furthermore, we optimise the algorithm using information from the internal structure of the clusters of accreted stars. We show that our approach yields good results in terms of both purity and completeness of clusters for galaxies with different types of accretion histories. We also evaluate the decrease in efficiency due to contamination by in situ stars. While for accreted-only haloes the algorithm matches well the recovered clusters with the individual progenitors and is able to recover accretion events up to a redshift of accretion $z_{\rm acc}\sim 3$, for accreted + in situ haloes it can only identify the more recent accretion events ($z_{\rm acc}<1$). However, the purity of the identified clusters remains remarkably high even in this case. Our results suggest that HDBSCAN can efficiently identify accreted debris in Milky Way-type galaxies in realistic conditions, however, it requires careful optimization to provide valid results.
@@
pubyear: 2025pagerange: Optimized HDBSCAN clustering for reconstructing the merger history of the Milky Way: applications and limitationsOptimized HDBSCAN clustering for reconstructing the merger history of the Milky Way: applications and limitations
@@ 1 Introduction
The increased availability of high-quality astrometric data from Gaia (Gaia Collaboration et al., 2016, 2023) and complementary chemical abundances from spectroscopic surveys such as APOGEE (Majewski et al., 2017), Gaia-ESO (Gilmore et al., 2012; Randich et al., 2013), LAMOST (Zhao et al., 2012), H3 (Conroy et al., 2019), or GALAH (Buder et al., 2021, 2025) has allowed the discovery of several substructures in the Galaxy. The debris identified so far appear as distinct features in the chemodynamical space (e.g., Klement et al., 2009; Helmi et al., 2018; Belokurov et al., 2018; Koppelman et al., 2019; Kruijssen et al., 2019; Ji et al., 2020; Naidu et al., 2020; Horta et al., 2023; Malhan & Rix, 2024).
@@
However, the identification of debris remains challenging in the presence of the background population and uncertainties introduced by instrumental effects (Brown et al., 2005), and given that overdensities originating from multiple progenitors overlap in kinematical space and share similar chemical abundances (Jean-Baptiste et al., 2017; Orkney et al., 2023; Mori et al., 2024).
@@
Machine learning methods, and clustering algorithms in particular, are well-suited for identifying halo substructures, given the natural tendency of stars from disrupted progenitors to cluster in kinematical space or display similar chemical abundances, in correspondence to the properties of progenitor galaxies they originate from. Indeed, clustering methods have shown their usefulness by detecting several new stellar substructures in the Milky Way and confirming some of the existing ones (e.g., Koppelman et al., 2019; Myeong et al., 2019; Borsato et al., 2020; Necib et al., 2020; Re Fiorentin et al., 2015; Ruiz-Lara et al., 2022; Yuan et al., 2020; Dodd et al., 2023).
@@
However, clustering methods also have some limitations when applied to Galactic archaeology. One example is the family of partitioning algorithms, such as $k$-means (MacQueen, 1967), which aim to divide the data into distinct groups by minimizing the intra-cluster variance. Their requirement to predefine the number of clusters makes them less ideal for identifying debris from Milky Way progenitors, where this number is not known a priori. These methods are more useful when the clusters correspond to known structures, such as multiple stellar populations of globular clusters (Fraix-Burnet & Davoust, 2015), or the distinct Galactic components (Mackereth et al., 2019). Moreover, the $k$-means algorithm is biased toward identifying spherically clustered groups of stars, which makes it unsuited for picking out tidal debris that typically displays an asymmetric distribution. For example, instead of “clumps” in the energy – angular momentum space, as expected from conservation of these two quantities for any progenitor, the overdensities in this space could be severely distorted as debris orbits in an evolving gravitational potential or interacts with other substructures in the host galaxy (e.g., Font et al., 2006; Gómez et al., 2010, 2013).
@@
Hierarchical clustering methods (see review of Murtagh & Contreras, 2012), in particular the so-called single-linkage methods, build a nested tree of clusters by progressively merging the two closest clusters in the parameter space (or feature space, in machine learning terminology) until every point belongs to the same cluster. Despite mimicking the hierarchical nature of galaxy formation, these methods incorporate every data point in a hierarchy which makes them sensitive to noise. Isolated points in the feature space can act as a bridge between clusters or distort the internal structure of clusters. Moreover, extracting a final, “flat” partition from the hierarchy is often not straightforward, as meaningful clusters, representing coherent debris from distinct accretion events, can appear at various levels within the hierarchy, reflecting differences in their spatial extent and detectability. For example, a tightly clustered debris from a recently accreted progenitor might remain distinct at a higher level in the hierarchy, whereas a more dispersed debris from an early accretion might merge with the background noise at a lower level. Using the single linkage-based clustering methodology developed by Lövdal et al. (2022) on the Auriga simulations, Thomas et al. (2025) found that only recently accreted (<67<6-7 Gyr ago) progenitors are efficiently identified. They also found that the method is prone to generate artificial detections, with a significant fraction of clusters being composed mainly of in situ stars.
@@
However, as with other clustering algorithms, the configuration of the clusters retrieved by HDBSCAN varies considerably with the choice of its internal parameters. An improper parameter selection can lead to either the fragmentation of larger, coherent structures or the erroneous aggregation of unrelated stars into spurious clumps.
In this study, we evaluate the effectiveness of HDBSCAN on simulated stellar haloes, where the “ground truth” is known (i.e., tidal debris is already identified and mapped to its corresponding progenitor galaxies, and contamination from in situ stars is also known). We use a sample of three Milky Way-mass galaxies from Auriga suite (Grand et al., 2017, 2024), which has been shown to form realistic disk-dominated galaxies similar to the Milky Way (Monachesi et al., 2019). The three galaxies are chosen so that they have significantly different assembly histories, in order to evaluate the variations in the performance of HDBSCAN with different chemodynamical distributions of tidal debris. We also incorporate a wider range of chemodynamical properties (features) in the analysis than previously used in similar clustering algorithms.
@@
Furthermore, the stellar halo of the Milky Way is known to have two distinct components, accreted and formed in situ (Carollo et al., 2007; Beers et al., 2012; Belokurov & Kravtsov, 2022). Cosmological simulations naturally produce dual haloes (Zolotov et al., 2009; Font et al., 2011; McCarthy et al., 2012; Tissera et al., 2013; Cooper et al., 2015). Because the two halo populations can have overlapping chemodynamical properties (Haywood et al., 2018; Di Matteo et al., 2019; Mori et al., 2024), the performance of the clustering algorithm generally decreases in the presence of in situ stars or can lead to the identification of spurious substructures. To rigorously validate the efficacy of our optimized HDBSCAN methodology, we first apply it to accreted-only components. This allows us to quantify the intrinsic ability of the algorithm in recovering accreted structures of varying morphologies and densities without the contamination from the in situ population, hence establishing a baseline for interpreting the results. With this baseline established, we then apply the algorithm to the fully mixed stellar halo.
@@ 2 The Auriga Simulations
The Auriga galaxy formation model incorporates a wide range of subgrid physics, including primordial and metal line cooling with self-shielding corrections against a spatially uniform, redshift-dependent UV background field (Faucher-Giguère et al., 2009) that completes reionization at $z\sim 6$. The interstellar medium (ISM) is treated using the two-phase model of Springel & Hernquist (2003), with stochastic star formation occurring in gas denser than $n=0.13\,\mathrm{cm}^{-3}$ assuming a Chabrier (2003) initial mass function. The model also accounts for stellar evolution, tracking the associated mass loss and chemical enrichment of the surrounding gas from asymptotic giant branch stars, Type Ia supernovae and Type II supernovae. Feedback is implemented through stellar wind particles released from Type II supernovae, as well as with radiative and thermal energy outbursts from active galactic nuclei with both radio and quasar modes. The model also accounts for the seeding and growth of supermassive black holes and includes a uniform primordial magnetic field of $10^{-14}\,\mathrm{G}$ (comoving).
@@
In our analysis, a progenitor of a Milky Way-mass galaxy is defined as a group of accreted stars [^3] We note that the term “stars” in the context of simulations refers to star particles, which are essentially single stellar populations. that share the same PeakMassIndex; henceforth, we will refer to these progenitors based on their respective numerical index, i.e., “prog. PeakMassIndex”. We motivate this definition by the fact that subhaloes tend to reach their maximum stellar mass just before accretion onto their host, that is, when they first cross the $R_{200}$ of their host. At around this time, which we denote $\tau_{\rm infall}$, most infalling satellites experience tidal and ram-pressure stripping, which results in a sharp decrease in their star formation (e.g., Kawata & Mulchaey 2008; Simpson et al. 2018; Font et al. 2022). We therefore evaluate the performance of the HDBSCAN algorithm by comparing the clusters of accreted stars identified by the algorithm with the PeakMassIndex-defined progenitor groups.
@@ 2.1 The simulations sample
The three host galaxies display significantly different present-day distributions of accreted stars, in either kinematics, chemical abundances or ages. Au 7 has a predominantly early formation epoch, and consequently the accreted debris is mostly phase-mixed at at $z=0$; in contrast, Au 25 has a late formation, and its debris is mostly in the form of cold stellar streams; Au27 has an assembly history more similar to the Milky Way, including an early major merger similar to the Gaia Enceladus-Sausage. Au 27 also has a large stellar disc and a bar at $z=0$.
In addition to the information from the merger trees, we also use the morphological classification of tidal debris in this Auriga suite performed by Riley et al. (2024). This classification groups accretion events into three categories, “intact”, “phase-mixed” and “stellar streams”, based on their configuration at $z=0$. An accretion event is classified as intact if the majority ($97\%$) of accreted stars are still bound to the system at present. The debris from disrupted satellites is considered either phase-mixed or as (cold) stellar stream based on a mass-dependent cut in the velocity dispersion, with the phase-mixed structures exhibiting a higher velocity dispersion than the streams.
For each galaxy, we consider only the accreted stars that formed in the main halo, defined as those accretion events that contribute at least 0.1% to the total number of accreted particles at $z=0$. This corresponds to a cutoff in the number of particles per progenitor ranging between 100 and 1000, depending on the system. We also exclude particles that are bound to surviving satellites and those from merger events labeled as “intact”, under the assumption that the identification of these systems would not require the application of clustering techniques.
@@ 2.1.3 Au27
The accreted halo in Au27 is dominated by stars originating from a massive satellite (stellar mass of $\sim 10^{9.5}~\mathrm{M_{\sun}}$), prog. 7219, accreted on a radial orbit at $z_{\rm acc}\sim 1.5$, similar to a Gaia Enceladus-Sausage event (Helmi et al., 2018; Belokurov et al., 2018). This contributed $\sim 40\%$ of the accreted star particles bound to the main halo at $z=0$. Two other significant, slightly less massive satellites were accreted just before, contributing 19% and, respectively, 14% to the mass of the accreted halo. Although most of the accreted debris are completely mixed in with in situ stars, 13% of the accreted stars still present stellar stream-like configurations; such an example is the remnant of the massive satellite, prog. 205, accreted $\sim 7~\mathrm{Gyr}$ ago. Au27 also presents a companion of similar mass as the Large Magellanic Cloud at $z=0$, which, however, is not considered in the analysis.
@@ 3.1 Optimizing the HDBSCAN parameters with OPTUNA
    the V-measure (Rosenberg & Hirschberg, 2007), an external entropy-based metric that compares the particle IDs in the HDBSCAN clusters to the ones in the progenitor groups defined by PeakMassIndex. It is defined by a combination of two criteria: homogeneity, which is satisfied if each cluster contains only particles of a single progenitor, and completeness, which is satisfied if all members of a given progenitor are assigned to the same cluster. The V-measure score ranges from 0 to 1, with higher scores indicating better clustering results. A score of 1 represents a perfectly homogeneous and complete clustering, meaning each cluster contains only particles from a single progenitor, and all particles of a given progenitor are grouped together in one cluster. A score of 0 indicates a completely random or poor clustering, where there is no relationship between the clusters and the true progenitor groups.
@@ 3.3 The 12-dimensional parameter (feature) space
We apply the HDBSCAN algorithm on a multidimensional parameter space comprised of ages, chemical abundance ratios and kinematics of halo stars. The kinematics are computed in a Cartesian coordinate system centred on the centre-of-mass of the host galaxy, and the $z$-axis aligned with the total angular momentum vector of the disc component. Two of the kinematical parameters are the integrals of motion (IoM): the energy $E$ and $L_{z}$, the component of the angular momentum along the $z$-axis, which is perpendicular to the plane of the stellar disc. We also include the perpendicular component of the angular momentum, which is the component in the xyx-y plane, $L_{\perp}=\sqrt{L_{x}^{2}+L_{y}^{2}}$. This is often used, even though it is known not to be fully conserved (Helmi et al., 1999). To these, we add the Cartesian positions and velocities ($x,y,z,v_{x},v_{y},v_{z}$). For chemical abundances, we use the $\alpha$-to-iron and iron-to-hydrogen chemical abundance ratios ($\alpha/\mathrm{Fe}]$, $\mathrm{Fe/H}]$), while for ages we use the lookback time since formation of star particles ($\tau_{\rm form}$). In summary, we choose a 12 dimensional (12-D) parameter space comprised of:
@@
Each of these stellar properties provides complementary information in the identification of progenitors. The IoMs are conserved quantities over time, which makes them ideal for identifying early mergers that appear as phase-mixed substructures at $z=0$ (Helmi & White, 1999). The phase-space coordinates are most effective for recently accreted debris, such as stellar streams, where stars still share similar positions and velocities (Johnston et al., 2008). Chemical abundance ratios are excellent tracers of the common chemical evolution of stars within satellite galaxies prior to accretion onto their host (Freeman & Bland-Hawthorn, 2002), while stellar ages allow the reconstruction of the star formation history from a single progenitor (Montalbán et al., 2021; Gallart et al., 2024).

### v2

@@ Abstract
Clustering algorithms can help reconstruct the assembly history of the Milky Way by identifying groups of stars sharing similar properties in a kinematical or chemical abundance space. Despite being promising tools, their efficiency has not yet been fully tested in a realistic cosmological framework. We investigate the effectiveness of the HDBSCAN clustering algorithm in the recovery of the progenitors of Milky Way-type galaxies, using several systems from the Auriga suite of simulations. We develop a methodology aimed at improving the efficiency of the algorithm and avoiding fragmentation: First, we use a 12-dimensional feature space including a range of chemodynamical properties and stellar ages; furthermore, we optimise the algorithm using information from the internal structure of the clusters of accreted stars. We show that our approach yields good results in terms of both purity and completeness of clusters for galaxies with different types of accretion histories. We also evaluate the decrease in efficiency due to contamination by in situ stars. While for accreted-only haloes the algorithm matches well the recovered clusters with the individual progenitors and is able to recover accretion events up to a redshift of accretion $z_{\rm acc}\sim 3$, for accreted + in situ haloes it can only identify the more recent accretion events ($z_{\rm acc}<1$). However, the purity of the identified clusters remains remarkably high even in this case. Our results suggest that HDBSCAN can efficiently identify accreted debris in Milky Way-type galaxies in realistic conditions, however, it requires careful optimization to provide valid results.
@@
pubyear: 2026pagerange: Optimized HDBSCAN clustering for reconstructing the merger history of the Milky Way: applications and limitationsOptimized HDBSCAN clustering for reconstructing the merger history of the Milky Way: applications and limitations
@@ 1 Introduction
The increased availability of high-quality astrometric data from Gaia (Gaia Collaboration et al., 2016, 2023) and complementary chemical abundances from spectroscopic surveys such as APOGEE (Majewski et al., 2017), Gaia-ESO (Gilmore et al., 2012; Randich et al., 2013), LAMOST (Zhao et al., 2012), H3 (Conroy et al., 2019), or GALAH (Buder et al., 2021, 2025) has allowed the discovery of several substructures in the Galaxy. The debris identified so far appear as distinct features in the chemodynamical space (e.g., Klement et al., 2009; Helmi et al., 2018; Belokurov et al., 2018; Koppelman et al., 2019; Kruijssen et al., 2019; Ji et al., 2020; Naidu et al., 2020; Horta et al., 2023; Malhan and Rix, 2024).
@@
However, the identification of debris remains challenging in the presence of the background population and uncertainties introduced by instrumental effects (Brown et al., 2005), and given that overdensities originating from multiple progenitors overlap in kinematical space and share similar chemical abundances (Jean-Baptiste et al., 2017; Orkney et al., 2023; Carrillo et al., 2024; Mori et al., 2024; Kizhuprakkat et al., 2024).
@@
Machine learning methods, and clustering algorithms in particular, are well-suited for identifying halo substructures, given the natural tendency of stars from disrupted progenitors to cluster in kinematical space or display similar chemical abundances, in correspondence to the properties of progenitor galaxies they originate from. Indeed, clustering methods have shown their usefulness by detecting several new stellar substructures in the Milky Way and confirming some of the existing ones (e.g., Koppelman et al., 2019; Myeong et al., 2019; Borsato et al., 2020; Necib et al., 2020; Re Fiorentin et al., 2015; Ruiz-Lara et al., 2022; Yuan et al., 2020; Dodd et al., 2023; Kim et al., 2025).
@@
However, clustering methods also have some limitations when applied to Galactic archaeology. One example is the family of partitioning algorithms, such as $k$-means (MacQueen, 1967), which aim to divide the data into distinct groups by minimizing the intra-cluster variance. Their requirement to predefine the number of clusters makes them less ideal for identifying debris from Milky Way progenitors, where this number is not known a priori. These methods are more useful when the clusters correspond to known structures, such as multiple stellar populations of globular clusters (Fraix-Burnet and Davoust, 2015), or the distinct Galactic components (Mackereth et al., 2019). Moreover, the $k$-means algorithm is biased toward identifying spherically clustered groups of stars, which makes it unsuited for picking out tidal debris that typically displays an asymmetric distribution. For example, instead of “clumps” in the energy – angular momentum space, as expected from conservation of these two quantities for any progenitor, the overdensities in this space could be severely distorted as debris orbits in an evolving gravitational potential or interacts with other substructures in the host galaxy (e.g., Font et al., 2006; Gómez et al., 2010, 2013).
@@
Hierarchical clustering methods (see review of Murtagh and Contreras, 2012), in particular the so-called single-linkage methods, build a nested tree of clusters by progressively merging the two closest clusters in the parameter space (or feature space, in machine learning terminology) until every point belongs to the same cluster. Despite mimicking the hierarchical nature of galaxy formation, these methods incorporate every data point in a hierarchy which makes them sensitive to noise. Isolated points in the feature space can act as a bridge between clusters or distort the internal structure of clusters. Moreover, extracting a final, “flat” partition from the hierarchy is often not straightforward, as meaningful clusters, representing coherent debris from distinct accretion events, can appear at various levels within the hierarchy, reflecting differences in their spatial extent and detectability. For example, a tightly clustered debris from a recently accreted progenitor might remain distinct at a higher level in the hierarchy, whereas a more dispersed debris from an early accretion might merge with the background noise at a lower level. Using the single linkage-based clustering methodology developed by Lövdal et al. (2022) on the Auriga simulations, Thomas et al. (2025) found that only recently accreted (<67<6-7 Gyr ago) progenitors are efficiently identified. They also found that the method is prone to generate artificial detections, with a significant fraction of clusters being composed mainly of in situ stars.
@@
However, as with other clustering algorithms, the configuration of the clusters retrieved by HDBSCAN varies considerably with the choice of its internal parameters. An improper parameter selection can lead to either the fragmentation of larger, coherent structures or the erroneous aggregation of unrelated stars into spurious clumps (Kim et al., 2025).
In this study, we evaluate the effectiveness of HDBSCAN on simulated stellar haloes, where the “ground truth” is known (i.e., tidal debris is already identified and mapped to its corresponding progenitor galaxies, and contamination from in situ stars is also known). We use a sample of three Milky Way-mass galaxies from Auriga suite (Grand et al., 2017, 2024), which has been shown to form realistic disk-dominated galaxies similar to the Milky Way (Grand et al., 2017; Orkney et al., 2022). The three galaxies are chosen so that they have significantly different assembly histories, in order to evaluate the variations in the performance of HDBSCAN with different chemodynamical distributions of tidal debris. We also incorporate a wider range of chemodynamical properties (features) in the analysis than previously used in similar clustering algorithms.
@@
Furthermore, the stellar halo of the Milky Way is known to have two distinct components, accreted and formed in situ (Carollo et al., 2007; Beers et al., 2012; Belokurov and Kravtsov, 2022). Cosmological simulations naturally produce dual haloes (Zolotov et al., 2009; Font et al., 2011; McCarthy et al., 2012; Tissera et al., 2013; Cooper et al., 2015). Because the two halo populations can have overlapping chemodynamical properties (Haywood et al., 2018; Di Matteo et al., 2019; Mori et al., 2024), the performance of the clustering algorithm generally decreases in the presence of in situ stars or can lead to the identification of spurious substructures. To rigorously validate the efficacy of our optimized HDBSCAN methodology, we first apply it to accreted-only components. This allows us to quantify the intrinsic ability of the algorithm in recovering accreted structures of varying morphologies and densities without the contamination from the in situ population, hence establishing a baseline for interpreting the results. With this baseline established, we then apply the algorithm to the fully mixed stellar halo.
@@ 2 The Auriga Simulations
The Auriga galaxy formation model incorporates a wide range of subgrid physics, including primordial and metal line cooling with self-shielding corrections against a spatially uniform, redshift-dependent UV background field (Faucher-Giguère et al., 2009) that completes reionization at $z\sim 6$ (Vogelsberger et al., 2013). The interstellar medium (ISM) is treated using the two-phase model of Springel and Hernquist (2003), with stochastic star formation occurring in gas denser than $n=0.13\,\mathrm{cm}^{-3}$ assuming a Chabrier (2003) initial mass function. The model also accounts for stellar evolution, tracking the associated mass loss and chemical enrichment of the surrounding gas from asymptotic giant branch stars (Karakas, 2010), Type Ia supernovae (Thielemann et al., 2003; Travaglio et al., 2004) and Type II supernovae (Portinari et al., 1998). Feedback is implemented through stellar wind particles released from Type II supernovae, as well as with radiative and thermal energy outbursts from active galactic nuclei with both radio and quasar modes. The model also accounts for the seeding and growth of supermassive black holes (Springel, 2005) and includes a uniform, comoving primordial magnetic field of $10^{-14}\,\mathrm{G}$ (Pakmor and Springel, 2013).
@@
In our analysis, a progenitor of a Milky Way-mass galaxy is defined as a group of accreted stars [^3] We note that the term “stars” in the context of simulations refers to star particles, which are essentially single stellar populations. that share the same PeakMassIndex; henceforth, we will refer to these progenitors based on their respective numerical index, i.e., “prog. PeakMassIndex”. We motivate this definition by the fact that subhaloes tend to reach their maximum stellar mass just before accretion onto their host, that is, when they first cross the $R_{200}$ of their host. At around this time, which we denote $\tau_{\rm infall}$, most infalling satellites experience tidal and ram-pressure stripping, which results in a sharp decrease in their star formation (e.g., Kawata and Mulchaey 2008; Simpson et al. 2018; Font et al. 2022). We therefore evaluate the performance of the HDBSCAN algorithm by comparing the clusters of accreted stars identified by the algorithm with the PeakMassIndex-defined progenitor groups.
@@ 2.1 The simulations sample
The three host galaxies display significantly different present-day distributions of accreted stars, in either kinematics, chemical abundances or ages. Au 7 has a predominantly early formation epoch, and consequently the accreted debris is mostly phase-mixed at $z=0$; in contrast, Au 25 has a late formation, and its debris is mostly in the form of cold stellar streams; Au27 has an assembly history more similar to the Milky Way, including an early major merger similar to the Gaia Enceladus-Sausage (GES) (Fattahi et al., 2019). Au 27 also has a large stellar disc and a bar at $z=0$.
In addition to the information from the merger trees, we also use the morphological classification of tidal debris in this Auriga suite performed by Riley et al. (2025) and Shipp et al. (2025). This classification groups accretion events into three categories, “intact”, “phase-mixed” and “stellar streams”, based on their configuration at $z=0$. An accretion event is classified as intact if the majority ($97\%$) of accreted stars are still bound to the system at present. The debris from disrupted satellites is considered either phase-mixed or as (cold) stellar stream based on a mass-dependent cut in the velocity dispersion, with the phase-mixed structures exhibiting a higher velocity dispersion than the streams.
For each galaxy, we consider only the accreted stars that contribute to the main halo, defined as those accretion events that contribute at least 0.1% to the total number of accreted particles at $z=0$. This corresponds to a cutoff in the number of particles per progenitor ranging between 100 and 1000, depending on the system. We also exclude particles that are bound to surviving satellites and those from merger events labeled as “intact”, under the assumption that the identification of these systems would not require the application of clustering techniques.
@@ 2.1.3 Au27
The accreted halo in Au27 is dominated by stars originating from a massive satellite (stellar mass of $\sim 10^{9.5}~\mathrm{M_{\sun}}$), prog. 7219, accreted on a radial orbit at $z_{\rm acc}\sim 1.5$, similar to a GES event (Helmi et al., 2018; Belokurov et al., 2018). This contributed $\sim 40\%$ of the accreted star particles bound to the main halo at $z=0$. Two other significant, slightly less massive satellites were accreted just before, contributing 19% and, respectively, 14% to the mass of the accreted halo. Although most of the accreted debris are completely mixed in with in situ stars, 13% of the accreted stars still present stellar stream-like configurations; such an example is the remnant of the massive satellite, prog. 205, accreted $\sim 7~\mathrm{Gyr}$ ago. Au27 also presents a companion of similar mass as the Large Magellanic Cloud at $z=0$, which, however, is not considered in the analysis.
@@ 3.1 Optimizing the HDBSCAN parameters with OPTUNA
    the V-measure (Rosenberg and Hirschberg, 2007), an external entropy-based metric that compares the particle IDs in the HDBSCAN clusters to the ones in the progenitor groups defined by PeakMassIndex. It is defined by a combination of two criteria: homogeneity, which is satisfied if each cluster contains only particles of a single progenitor, and completeness, which is satisfied if all members of a given progenitor are assigned to the same cluster. The V-measure score ranges from 0 to 1, with higher scores indicating better clustering results. A score of 1 represents a perfectly homogeneous and complete clustering, meaning each cluster contains only particles from a single progenitor, and all particles of a given progenitor are grouped together in one cluster. A score of 0 indicates a completely random or poor clustering, where there is no relationship between the clusters and the true progenitor groups.
@@ 3.3 The 12-dimensional parameter (feature) space
We apply the HDBSCAN algorithm on a multidimensional parameter space comprised of ages, chemical abundance ratios and kinematics of halo stars. The kinematics are computed in a Cartesian coordinate system centred on the centre-of-mass of the host galaxy, and the $z$-axis aligned with the total angular momentum vector of the disc component. Two of the kinematical parameters are the integrals of motion (IoM): the energy $E$ and $L_{z}$, the component of the angular momentum along the $z$-axis, which is perpendicular to the plane of the stellar disc. We also include the perpendicular component of the angular momentum, which is the component in the xyx-y plane, $L_{\perp}=\sqrt{L_{x}^{2}+L_{y}^{2}}$. This is often used, even though it is known to not be fully conserved (Helmi et al., 1999). To these, we add the Cartesian positions and velocities [^4] We experimented with the removal of positions and velocities from the clustering features, but we found a decrease in the performance of the model compared with including them. We therefore decided to incorporate positions and velocities among the features as they may add information (for example, the location of particles in the host galaxy, which otherwise might be hard to disentangle from the angular momentum features). ($x,y,z,v_{x},v_{y},v_{z}$). For chemical abundances, we use the $\alpha$-to-iron and iron-to-hydrogen chemical abundance ratios ($\alpha/\mathrm{Fe}]$, $\mathrm{Fe/H}]$), while for ages we use the lookback time since formation of star particles ($\tau_{\rm form}$). In summary, we choose a 12 dimensional (12-D) parameter space comprised of:
@@
Each of these stellar properties provides complementary information in the identification of progenitors. The IoMs are conserved quantities over time, which makes them ideal for identifying early mergers that appear as phase-mixed substructures at $z=0$ (Helmi and White, 1999). The phase-space coordinates are most effective for recently accreted debris or progenitors that are actively being disrupted, where the positions and velocities can be highly correlated (Johnston et al., 2008). Chemical abundance ratios are excellent tracers of the common chemical evolution of stars within satellite galaxies prior to accretion onto their host (Freeman and Bland-Hawthorn, 2002), while stellar ages allow the reconstruction of the star formation history from a single progenitor (Montalbán et al., 2021; Gallart et al., 2024).